OpenStudy (anonymous):

Which number is a solution of 5m<6m+6 a) -8 b) -5 c) -6 d) -7

OpenStudy (ahsome):

Is this it: \[5m<6m+6\] ?

OpenStudy (ahsome):

Assuming it is Equation: \(5m<6m+6\) First, take away both sides by \(6m\) to move it to the LHS \(5m<6m+6\) \(5m-6m<6m+6-6m\) \(5m-6m<\cancel{6m}+6\cancel{-6m}\) \(5m-6m<6\) \(-m<6\) Divide both sides by \(-1\). Remember, when you divide or multiply by a negative number in an inequality, you flip the signs \(-m<6\) \[\frac{-1\times m}{-1}>\frac{6}{-1}\]\[\frac{\cancel{-1}\times m}{\cancel{-1}}>\frac{6}{-1}\]\[m>\frac{6}{-1}\]\[m=-6\] TADA

Directrix (directrix):

@Ahsome The solution to your inequality is m > -6. So, -6 is not the answer to the posted question because -6 is not > -6.

OpenStudy (anonymous):

im so confused...

OpenStudy (ahsome):

Sorry, whoops

OpenStudy (ahsome):

I mean \(m\) HAS to be bigger than \(-6\) Which gives \(-5\) as the only answer

OpenStudy (ahsome):

Sorry @Directrix

OpenStudy (anonymous):

thanks guys! (:

OpenStudy (ahsome):

NP :)

Directrix (directrix):

@Ahsome I agree with your revised answer. It is easy to get mixed up on inequalities.

OpenStudy (ahsome):

Yes, yes it can be @Directrix